# Characterizing the Burst Error Correction Ability of Quantum Cyclic Codes

### v1

@@ VI Conclusion and Discussion
Regarding the future work, how to dertermine the burst error correction ability of QCCs with mathematical methods is quite useful. Although the time complexity of Algorithms 1&\&&2 is polynomial, the exhaustive searching complexity is high when the code length is relatively large. On the other hand, whether the quantum error-trapping decoder is degenerate quantum maximum likelihood decoding is unknown. It is an interesting problem to find degenerate quantum maximum likelihood decoding for quantum cyclic codes.
@@ Appendix A Proof of the Quantum Reiger Bound
If we append $\lfloor\frac{n}{2}\rfloor$ ancilla qubits $|0\cdots 0\rangle$ to the first sub-block, and append $n-\lfloor\frac{n}{2}\rfloor$ ancilla qubits $|0\cdots 0\rangle$ to the second sub-block, then the original encoded block has spawned two offspring, the first one with located burst errors of length at most $\lfloor\frac{n}{2}\rfloor$, and the second one with located burst errors of length at most $n-\lfloor\frac{n}{2}\rfloor$. If we were able to correct the two located burst errors in each of the offspring (see Lemma 11), we would obtain two identical copies of the parent encoded block, which is a contradiction with the quantum no-cloning theorem [1]. Therefor we must have $n>4\ell.$
@@
where $1\leq x<y\leq n,y-x+1=2\ell$, and $e_{i}(1\leq i\leq n)$ are Pauli matrices. The set $E(x,y)$ of burst errors to be corrected is the set of all Pauli operators, where each acts trivially on the qubits $1$ to $x-1$ and on the qubits $y+1$ to $n$ (except $x=1$ and $y=n$). Then each error in $E(x,y)$ has a burst length of at most $2\ell$. But now, for each $E_{a}$ and $E_{b}$ in $E(x,y)$, the product $E_{a}^{\dagger}E_{b}$ also has a burst of length at most $2\ell$. Therefore, the burst error-correcting criterion (2) is satisfied for all $E_{a,b}\in E$, provided $Q$ is an $\ell$ burst error correction code. ∎
@@
The proof follows closely by that of the quantum Singleton bound given by Preskill (see [57, p.32] and [1, p.568]).
@@ Appendix B Proof of Theorem 2
If item 1) holds, we know that ${C}$ can correct any burst error of length $\ell$ or less according to Lemma 4. If $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank for some $11\leq\imath\leq n-2\ell+$ 1, it means that a number of columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ are linearly dependent. It is known that every set of linearly dependent columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ corresponds to a pair of burst errors $(e,f)$ such that $He^{T}=Hf^{T}$. We define a set of all such burst errors as $\mathcal{E}_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$. We need to verify that whether such burst errors in $\mathcal{E}_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ are degenerate or not.
@@
According to Lemma 2, we can construct a ${Q}=[[n,2k-n]]$ QBECC which can correct any quantum burst error of length $\mathcal{L}$ or less. ∎
@@ Appendix C Proofs of Lemma 7, Lemma 9, and Theorem 3
Let $[\mu_{i}]=(u_{i1},\cdots,u_{im})$ for $1\leq i\leq n$. Then there is $[\mu]=([\mu_{1}],\cdots,[\mu_{n}])\in\mathbf{D}$. For arbitrary two codewords $\mu^{(1)}=(\mu_{1}^{(1)},\cdots,\mu_{n}^{(1)})\in C$ and $\mu^{(2)}=(\mu_{1}^{(2)},\cdots,\mu_{n}^{(2)})\in C$, if $\mu^{(1)}\neq\mu^{(2)}$, we have $\mu_{\mathbf{i}}^{(1)}\neq\mu_{\mathbf{i}}^{(2)}$ for some $1\leq\mathbf{i}\leq n$. Then there must be $[\mu^{(1)}]\neq[\mu^{(2)}]$.
@@
Let $\hat{B}_{\imath}=[\imath,\imath+\hbar-1]$, where $11\leq\imath\leq n-2\hbar-$ 1. We have
@@
where $A_{\imath+\hbar}$ is the (ı+)italic-$(\imath+\hbar)$th column of $\mathcal{M}^{(\hbar+1)}_{{B}_{\imath}}$, and $B_{r-\hbar}$ is the $(r-\hbar)$th row of $\mathcal{M}^{\hbar}_{\hat{B}_{\imath}}$. Since RS codes saturate the Reiger bound, $\mathcal{M}^{\hbar}_{\hat{B}_{\imath}}$ is of full rank and is equal to $\hbar$. Then $\mathcal{M}^{(\hbar+1)}_{\hat{B}_{\imath}}$ must be greater than or equal to $\hbar-1$. Therefore we have $\hbar-1\leq rank(\mathcal{M}^{(\hbar+1)}_{B_{\imath}})\leq\hbar$. ∎
@@
Let ${D}=[{C}]$ be the $q$-ary expansion of ${C}$ under the self-dual basis. Then there is ${D}^{\bot}\subseteq{D}$ according to Lemma 6. Moreover, ${D}$ can correct any quantum burst error of length $(\hbar-1)m+1$ or less according to Lemma 8. Let $\widetilde{e}$ and $\widetilde{f}$ be arbitrary two burst error over $\mathbb{F}_{q}$ such that
$(\hbar-1)m+1\leq\textrm{bl}(\widetilde{e})\leq\mathcal{L},\textrm{ and }1\leq\textrm{bl}(\widetilde{f})\leq\mathcal{L}.$ (30)
@@
Suppose that $\widetilde{e}+\widetilde{f}\in{D}\setminus{D}^{\bot}$, then $e+f\in{C}\setminus{C}^{\bot}$ according to Lemma 7. It is a contradiction with Eq. (11). Then there exists a $Q=[[n,2k-n]]$ quantum RS code which can correct any quantum burst error of length $\mathcal{L}$ or less according to Lemma 1 and Lemma 6.

### v2

@@ VI Conclusion and Discussion
Regarding the future work, how to determine the burst error correction ability of QCCs with mathematical methods is quite useful. Although the time complexity of Algorithms 1&\&2 is polynomial, the exhaustive searching complexity is high when the code length is relatively large. On the other hand, whether the quantum error-trapping decoder is degenerate, quantum maximum likelihood decoding is unknown. It is an interesting problem to find degenerate quantum maximum likelihood decoding for quantum cyclic codes.
Acknowledgment
The authors would like to thank the Editor and the anonymous referees for their valuable comments and suggestions for improving the presentation of their article.
@@ Appendix A Proof of the Quantum Reiger Bound
If we append $\lfloor\frac{n}{2}\rfloor$ ancilla qubits $|0\cdots 0\rangle$ to the first sub-block, and append $n-\lfloor\frac{n}{2}\rfloor$ ancilla qubits $|0\cdots 0\rangle$ to the second sub-block, then the original encoded block has spawned two offspring, the first one with located burst errors of length at most $\lfloor\frac{n}{2}\rfloor$, and the second one with located burst errors of length at most $n-\lfloor\frac{n}{2}\rfloor$. If we were able to correct the two located burst errors in each of the offspring (see Lemma 11), we would obtain two identical copies of the parent encoded block, which is a contradiction with the quantum no-cloning theorem [41]. Therefore we must have $n>4\ell.$
@@
where $1\leq x<y\leq n,y-x+1=2\ell$, and $e_{i}(1\leq i\leq n)$ are Pauli matrices. The set $E(x,y)$ of burst errors to be corrected is the set of all Pauli operators, where each acts trivially on the qubits $1$ to x1x-1 and on the qubits $y+1$ to $n$ (except $x=1$ and $y=n$). Then each error in $E(x,y)$ has a burst length of at most $2\ell$. But now, for each $E_{a}$ and $E_{b}$ in $E(x,y)$, the product $E_{a}^{\dagger}E_{b}$ also has a burst of length at most $2\ell$. Therefore, the burst error-correcting criterion (2) is satisfied for all $E_{a,b}\in E$, provided $Q$ is an $\ell$ burst error correction code. ∎
@@
The proof follows closely by that of the quantum Singleton bound given by Preskill (see [44, p.32] and [41, p.568]).
@@ Appendix B Proof of Theorem 2
If item 1) holds, we know that C{C} can correct any burst error of length $\ell$ or less according to Lemma 4. If $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank for some $1\leq\imath\leq n-2\ell+1$, it means that a number of columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ are linearly dependent. It is known that every set of linearly dependent columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ corresponds to a pair of burst errors $(e,f)$ such that $He^{T}=Hf^{T}$. We define a set of all such burst errors as $\mathcal{E}_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$. We need to verify that whether such burst errors in $\mathcal{E}_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ are degenerate or not.
@@
According to Lemma 2, we can construct a Q=[[n,2kn]]{Q}=[[n,2k-n]] QBECC which can correct any quantum burst error of length $\mathcal{L}$ or less.
@@ Appendix C Proofs of Lemma 7, Lemma 9, and Theorem 3
Let $\mu_{i}]=(u_{i1},\cdots,u_{im})$ for $1\leq i\leq n$. Then there is $\mu]=([\mu_{1}],\cdots,[\mu_{n}])\in\mathbf{D}$. For arbitrary two codewords $\mu^{(1)}=(\mu_{1}^{(1)},\cdots,\mu_{n}^{(1)})\in C$ and $\mu^{(2)}=(\mu_{1}^{(2)},\cdots,\mu_{n}^{(2)})\in C$, if $\mu^{(1)}\neq\mu^{(2)}$, we have $\mu_{\mathbf{i}}^{(1)}\neq\mu_{\mathbf{i}}^{(2)}$ for some $1\leq\mathbf{i}\leq n$. Then there must be $\mu^{(1)}]\neq[\mu^{(2)}]$.
@@
Let $\hat{B}_{\imath}=[\imath,\imath+\hbar-1]$, where $1\leq\imath\leq n-2\hbar-1$. We have
@@
where $A_{\imath+\hbar}$ is the $(\imath+\hbar)$th column of $\mathcal{M}^{(\hbar+1)}_{{B}_{\imath}}$, and $B_{r-\hbar}$ is the $(r-\hbar)$th row of $\mathcal{M}^{\hbar}_{\hat{B}_{\imath}}$. Since RS codes saturate the Reiger bound, $\mathcal{M}^{\hbar}_{\hat{B}_{\imath}}$ is of full rank and is equal to $\hbar$. Then $\mathcal{M}^{(\hbar+1)}_{\hat{B}_{\imath}}$ must be greater than or equal to $\hbar-1$. Therefore we have $\hbar-1\leq rank(\mathcal{M}^{(\hbar+1)}_{B_{\imath}})\leq\hbar$. ∎
@@
Let D=[C]{D}=[{C}] be the $q$-ary expansion of C{C} under the self-dual basis. Then there is ${D}^{\bot}\subseteq{D}$ according to Lemma 6. Moreover, D{D} can correct any quantum burst error of length $(\hbar-1)m+1$ or less according to Lemma 8. Let $\widetilde{e}$ and $\widetilde{f}$ be arbitrary two burst error over $\mathbb{F}_{q}$ such that
[…]
@@
Suppose that $\widetilde{e}+\widetilde{f}\in{D}\setminus{D}^{\bot}$, then $e+f\in{C}\setminus{C}^{\bot}$ according to Lemma 7. It is a contradiction with Eq. (11). Then there exists a Q=[[n,2kn]]Q=[[n,2k-n]] quantum RS code which can correct any quantum burst error of length $\mathcal{L}$ or less according to Lemma 1 and Lemma 6.
Jihao Fan received the B.S. degree in mathematics from Lanzhou University, Lanzhou, China, in 2009, and the Ph.D. degree in computer software and theory from Southeast University, Nanjing, China, in 2016. He was a joint-training Ph.D student at Sydney University, Australia from 2014 to 2015. He is currently an Associate Professor with the Nanjing University of Science and Technology, Nanjing, China. His research interests include classical and quantum coding theory, information theory, and code-based masking.
Min-Hsiu Hsieh has been the director of the Hon Hai Quantum Computing Research Center in Taiwan since January 2021. Prior to this role, he was an Associate Professor at the University of Technology Sydney in Australia and a member of the UTS Centre for Quantum Software and Information from 2014 to 2020. From 2010 to 2012, he was a postdoctoral research fellow at the University of Cambridge in the UK, where he was affiliated with the Centre for Quantum Information and Foundations. Before that, he worked as a research scientist with the ERATO-SORST Quantum Computation and Information Project in Japan, which was funded by the Japan Science and Technology Agency. During this time, he was also affiliated with the Department of Computer Science at the University of Tokyo. From 2014 to 2018, he held an Australian Research Council Future Fellowship. His research interests include general topics in quantum information and computation.